OpenStudy (anonymous):

If y varies directly with x. x is 30 y is 10 what is x when y is 9?

OpenStudy (jhannybean):

y varies directly with x is written as \(\large y=cx\). c is a constant . Now did you mean "x is 30 when y =10?

OpenStudy (anonymous):

The problem just says x is 30 when y is 1...

OpenStudy (jhannybean):

Ok.

OpenStudy (jhannybean):

So \(10=30c\) we solve for constant c to create another relationship when y =9. c= 1/3. So now when y= 9, we substitute c back into our original formula \(\large \color{red} {y=cx}\) but this time we have y = 9,c=1/3. 9=1/3 x where x = 9(3) = 27.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (jhannybean):

no problem :)